# Towards Principled Evaluations of Sparse Autoencoders for Interpretability and Control

### v1

@@ 1 Introduction
  • We apply this method to the IOI task, and show these dictionaries exhibit three desirable properties in the task’s context: (1) sufficiency/necessity of activation reconstructions, (2) sparse controllability of model behavior via feature editing, and (3) interpretability of the features consistent with their causal role111 As a by-product, our supervised feature dictionaries also demonstrate that activations for the task can be usefully disentangled in a way that adheres to the linear representation hypothesis and exhibits superposition. This is the first time such a disentanglement has been achieved in a realistic LLM task to the best of our knowledge..
@@ 3.1 Overview and Motivation
Step 2: Compute high-quality supervised dictionaries using the attributes. The second step is to learn and validate supervised dictionaries computed using the attribute labels $a_{i}(p),p\sim\mathcal{D}$ (methodology described in Section 4). In particular, not all attribute configurations will result in dictionaries with good approximation, control and interpretability for the task222 To get high-quality supervised dictionaries, we should choose attributes that completely capture the task-relevant information in the prompt, and are moreover compatible with the intermediate states of the model’s computation on the task. Validating that our attributes satisfy this is a non-trivial but key prerequisite for our evaluation to be meaningful. We do this for the IOI task in Section 4, where we show the supervised dictionaries score highly on all tests described in the current section. .
@@ 3.2 Test 1: Sufficiency and Necessity of Dictionary Reconstructions for the Task
Our first test checks if activation reconstructions $\widehat{\mathbf{a}}$ as a whole are sufficient and necessary for the model to perform the task [^3] This test is analogous to the faithfulness/completeness test from Wang et al. (2023); the sufficiency test is also widely used in the literature to evaluate feature dictionaries.. To evaluate sufficiency, we intervene by replacing internal activations $\mathbf{a}$ with their reconstructions $\widehat{\mathbf{a}}$, and measure the drop in performance on the task. To evaluate necessity, we intervene by replacing activations $\mathbf{a}$ with $\mathbb{E}_{\mathcal{D}}\left[\mathbf{a}\right]+\left(\mathbf{a}-\widehat{\mathbf{a}}\right)$, and compare the resulting drop in performance to the effect of replacing activations with $\mathbb{E}_{\mathcal{D}}\left[\mathbf{a}\right]$ alone (also known as mean ablation; see Appendix A.2 for motivation) 444 This test has very few assumptions: it is independent of the attributes we have chosen to describe inputs with, and of any interpretation assigned to the features. However, it only measures the quality of the dictionary as a whole, and not how effectively and sparsely it disentangles any hypothetical concepts in the activations. We include this test because poor performance indicates a fundamental failure of the feature dictionary relative to the task. .
@@ 3.3 Test 2: Sparse Controllability of Attributes
With this test, we want to measure the degree to which the feature dictionary can be used to control the model’s behavior on the task by editing intermediate representations of attribute values. To use a feature dictionary for editing, we write an activation as a linear combination of dictionary vectors, and try to remove/add dictionary elements to the combination in order to achieve the desired change. To evaluate such an intervention, we consider both (1) the number of features in the dictionary that need to be changed in order to achieve a given change in outputs (editing is trivial if e.g. activations are dense sums of random features and we change many of them), and (2) the geometric magnitude of the edit (editing is trivial if you are willing to throw away the entire activation vector and replace it with the target). Jump ahead to Figure 4 to see what Pareto curves for these parameters look like in our experiments.
@@
Implementation design space. To achieve independence of interpretations, we frame the problem as a combinatorial optimization problem over the feature dictionary: given an activation $\mathbf{a}_{s}$ for prompt $p_{s}$ and a counterfactual activation $\mathbf{a}_{t}$ for a prompt $p_{t}$ which differs from $p_{s}$ only in the values of the attributes we wish to edit, we can optimize over subsets of features active in $\widehat{\mathbf{a}_{s}}$ to subtract from $\mathbf{a}_{s}$, and subsets of features active in $\widehat{\mathbf{a}_{t}}$ to add to $\mathbf{a}_{s}$. There are multiple optimization objectives possible, such as making the edited activation similar to $\mathbf{a}_{t}$ geometrically, or making model behavior on the edited activation similar to model behavior on $\mathbf{a}_{t}$. The optimization may be constrained by the number of features changed and/or the magnitude of edits 666 There are many possible ways to instantiate such an optimization; in Section 5 we present one such way based on a greedy algorithm minimizing $\ell_{2}$ distance in activation space. Further methodological notes, such as how we implement edits in multiple model components at once, are given in Appendix A.2. . Importantly, this optimization proceeds on a per-prompt basis, which means that it can select different features to edit depending on the prompt being edited. This means it will not disadvantage dictionaries that do not dedicate a uniform w.r.t. all prompts set of features to each attribute (which would be a major assumption to impose).
@@ 4.1 Computing Supervised Feature Dictionaries
Choosing attributes for the IOI task. Not every set of attributes will result in a good approximation of the model’s internal activations; in fact, we find that the choice of attributes is a crucial modeling decision. Recall that, according to Wang et al. (2023), each IOI prompt $p$ is described by three properties influencing how $p$ is processed in the IOI circuit: the subject (S) and indirect object (IO) names, and their relative position (Pos). Motivated by this, we chose the attributes S, IO, and Pos to describe each prompt. We experimented with other choices of attributes, but did not find them to be more successful in our tests (see Appendix A.8 for details999 In particular, we found that another set of attributes, though more expressive in principle, learns to approximate the features we get using the S, IO and Pos attributes through a change-of-variables-like transformation.). We emphasize that there are many other imaginable choices of attributes; see Appendix A.10 for further discussion.
@@ 5.1 Methodology
Figure 4: Pareto curves for attribute editing using task-specific SAEs for select cross-sections of the IOI circuit. The x-axis measures the weight removed by an edit (averaged over the attention heads in the cross-section), affine-transformed so that a value of 0 indicates the weight removed by the corresponding supervised edit, and a value of 1 indicates removing $\approx$ all of the mass of the features. The y-axis is an affine transform of the fraction of examples for which the edit results in the same next-token prediction as the ground-truth edit, with a value of 00 corresponding to no intervention, and a value of $1$ corresponding to the supervised edits. The thick lines use our interpretation-agnostic editing method, while the dashed lines choose which features to edit based on $F_{1}$ scores for the corresponding attribute. In both cases, we edit 2, 4 or 6 features (a higher magnitude score indicates editing more features).
@@ 5.2 Results
Causal evaluation: sparse control via interpretable features. Here, results are also moderately encouraging. We find that, for the task-specific SAEs, editing using the high-$F_{1}$-score features w.r.t. a given attribute as a guide performs not much worse than the interpretation-agnostic editing method. Results are provided in Figure 4 in the form of a Pareto comparison with interpretation-agnostic editing, as well as Appendix Figure 28. However, for full-distribution SAEs, we again need to edit a high number of features to achieve a noticeable effect (results in Appendix Figure 29).
@@ A.13 Additional observations on feature interpretations
Generalization of features to the full distribution. We sample prompts from OpenWebText and visualize prompts that highly activate a feature. We do this for the name mover queries and outputs. For the outputs, we calculate the direct-feature-attribution (DFA) metric that calculates per position how large the contribution of its values to activating the feature is. Thus, it informs what position led to the feature being activated.
@@ A.15 Feature over-splitting in a mixture of gaussians toy model
We implemented this using $10^{5}$ samples, with a grid of 100 values for $\gamma$ in $[0,5]$ and 20 values for $k$ in $[0,2]$, over a grid of 100 values of $\lambda$ in $[0,20]$. We verify that the best values chosen for each $\lambda$ are not on the edges of the grid; the resulting curve of best total loss values versus $\lambda$ is shown in Figure TODO (blue), cut off at $\lambda\approx 7$, beyond which the selected SAE activates for $<2\%$ of examples in the components of the mixture.

### v2

@@ Abstract
footnotetext: : Joint contribution. Correspondence to: aleksandar.makelov@gmail.com.
@@ 1 Introduction
  • We apply this method to the IOI task, and show these dictionaries exhibit three desirable properties in the task’s context: (1) sufficiency/necessity of activation reconstructions, (2) sparse controllability of model behavior via feature editing, and (3) interpretability of the features consistent with their causal role000 As a by-product, our supervised feature dictionaries also demonstrate that activations for the task can be usefully disentangled in a way that adheres to the linear representation hypothesis and exhibits superposition. This is the first time such a disentanglement has been achieved in a realistic LLM task to the best of our knowledge..
@@ 3.1 Overview and Motivation
Step 2: Compute high-quality supervised dictionaries using the attributes. The second step is to learn and validate supervised dictionaries computed using the attribute labels $a_{i}(p),p\sim\mathcal{D}$ (methodology described in Section 4). In particular, not all attribute configurations will result in dictionaries with good approximation, control and interpretability for the task111 To get high-quality supervised dictionaries, we should choose attributes that completely capture the task-relevant information in the prompt, and are moreover compatible with the intermediate states of the model’s computation on the task. Validating that our attributes satisfy this is a non-trivial but key prerequisite for our evaluation to be meaningful. We do this for the IOI task in Section 4, where we show the supervised dictionaries score highly on all tests described in the current section. .
@@ 3.2 Test 1: Sufficiency and Necessity of Dictionary Reconstructions for the Task
Our first test checks if activation reconstructions $\widehat{\mathbf{a}}$ as a whole are sufficient and necessary for the model to perform the task [^2] This test is analogous to the faithfulness/completeness test from Wang et al. (2023); the sufficiency test is also widely used in the literature to evaluate feature dictionaries.. To evaluate sufficiency, we intervene by replacing internal activations $\mathbf{a}$ with their reconstructions $\widehat{\mathbf{a}}$, and measure the drop in performance on the task. To evaluate necessity, we intervene by replacing activations $\mathbf{a}$ with $\mathbb{E}_{\mathcal{D}}\left[\mathbf{a}\right]+\left(\mathbf{a}-\widehat{\mathbf{a}}\right)$, and compare the resulting drop in performance to the effect of replacing activations with $\mathbb{E}_{\mathcal{D}}\left[\mathbf{a}\right]$ alone (also known as mean ablation; see Appendix A.2 for motivation) 333 This test has very few assumptions: it is independent of the attributes we have chosen to describe inputs with, and of any interpretation assigned to the features. However, it only measures the quality of the dictionary as a whole, and not how effectively and sparsely it disentangles any hypothetical concepts in the activations. We include this test because poor performance indicates a fundamental failure of the feature dictionary relative to the task. .
@@ 3.3 Test 2: Sparse Controllability of Attributes
With this test, we want to measure the degree to which the feature dictionary can be used to control the model’s behavior on the task by editing intermediate representations of attribute values. To use a feature dictionary for editing, we write an activation as a linear combination of dictionary vectors, and try to remove/add dictionary elements to the combination in order to achieve the desired change. To evaluate such an intervention, we consider both (1) the number of features in the dictionary that need to be changed in order to achieve a given change in outputs (editing is trivial if e.g. activations are dense sums of random features and we change many of them), and (2) the geometric magnitude of the edit (editing is trivial if you are willing to throw away the entire activation vector and replace it with the target). Jump ahead to Figure 4 to see what trade-offs for these parameters look like in our experiments.
@@
Implementation design space. To achieve independence of interpretations, we frame the problem as a combinatorial optimization problem over the feature dictionary: given an activation $\mathbf{a}_{s}$ for prompt $p_{s}$ and a counterfactual activation $\mathbf{a}_{t}$ for a prompt $p_{t}$ which differs from $p_{s}$ only in the values of the attributes we wish to edit, we can optimize over subsets of features active in $\widehat{\mathbf{a}_{s}}$ to subtract from $\mathbf{a}_{s}$, and subsets of features active in $\widehat{\mathbf{a}_{t}}$ to add to $\mathbf{a}_{s}$. There are multiple optimization objectives possible, such as making the edited activation similar to $\mathbf{a}_{t}$ geometrically, or making model behavior on the edited activation similar to model behavior on $\mathbf{a}_{t}$. The optimization may be constrained by the number of features changed and/or the magnitude of edits 555 There are many possible ways to instantiate such an optimization; in Section 5 we present one such way based on a greedy algorithm minimizing $\ell_{2}$ distance in activation space. Further methodological notes, such as how we implement edits in multiple model components at once, are given in Appendix A.2. . Importantly, this optimization proceeds on a per-prompt basis, which means that it can select different features to edit depending on the prompt being edited. This means it will not disadvantage dictionaries that do not dedicate a uniform w.r.t. all prompts set of features to each attribute (which would be a major assumption to impose).
@@ 4.1 Computing Supervised Feature Dictionaries
Choosing attributes for the IOI task. Not every set of attributes will result in a good approximation of the model’s internal activations; in fact, we find that the choice of attributes is a crucial modeling decision. Recall that, according to Wang et al. (2023), each IOI prompt $p$ is described by three properties influencing how $p$ is processed in the IOI circuit: the subject (S) and indirect object (IO) names, and their relative position (Pos). Motivated by this, we chose the attributes S, IO, and Pos to describe each prompt. We experimented with other choices of attributes, but did not find them to be more successful in our tests (see Appendix A.8 for details888 In particular, we found that another set of attributes, though more expressive in principle, learns to approximate the features we get using the S, IO and Pos attributes through a change-of-variables-like transformation.). We emphasize that there are many other imaginable choices of attributes; see Appendix A.10 for further discussion.
@@ 5.1 Methodology
Figure 4: Trade-offs between edit magnitude and edit success for attribute editing using task-specific SAEs for select cross-sections of the IOI circuit. The x-axis measures the weight removed by an edit (averaged over the attention heads in the cross-section), affine-transformed so that a value of 0 indicates the weight removed by the corresponding supervised edit, and a value of 1 indicates removing $\approx$ all of the mass of the features. The y-axis is an affine transform of the fraction of examples for which the edit results in the same next-token prediction as the ground-truth edit, with a value of 00 corresponding to no intervention, and a value of $1$ corresponding to the supervised edits. The thick lines use our interpretation-agnostic editing method, while the dashed lines choose which features to edit based on $F_{1}$ scores for the corresponding attribute. In both cases, we edit 2, 4 or 6 features (a higher magnitude score indicates editing more features).
@@ 5.2 Results
Causal evaluation: sparse control via interpretable features. Here, results are also moderately encouraging. We find that, for the task-specific SAEs, editing using the high-$F_{1}$-score features w.r.t. a given attribute as a guide performs not much worse than the interpretation-agnostic editing method. Results are provided in Figure 4 in the form of a comparison with interpretation-agnostic editing, as well as Appendix Figure 28. However, for full-distribution SAEs, we again need to edit a high number of features to achieve a noticeable effect (results in Appendix Figure 29).
@@ Acknowledgements
Author Contributions
AM designed and ran the algorithms for computing supervised feature dictionaries and observed their theoretical properties; trained task-specific SAEs; designed the tests for sufficiency/necessity of reconstructions, sparse controllability and causal faithfulness of feature descriptions; ran these tests on full-distribution and task-specific SAEs; observed the occlusion/oversplitting phenomena in task-specific SAEs and developed the toy models for them; wrote all sections of the paper unless otherwise specified.
GL trained full-distribution SAEs; designed and ran the automatic feature scoring algorithms for them, and ran most of the interpretability evaluations for full-distribution SAEs; contributed to the sections on interpretability of full-distribution SAE features, created Figures 1 and 5.
NN provided guidance on research goals and prioritization, as well as writing, throughout the entire project; proposed the initial idea of finding supervised feature dictionaries for the IOI task, and using them to evaluate (unsupervised) feature dictionaries; refined the search for algorithms to compute the supervised dictionaries.
@@ A.13 Additional observations on feature interpretations
Generalization of features to the full distribution. We sample prompts from OpenWebText and visualize prompts that highly activate a feature. We do this for the name mover queries and outputs. For the outputs, we calculate the direct-feature-attribution (DFA) metric first proposed by Krzyzanowski et al. (2024) that calculates per position how large the contribution of its values to activating the feature is. Thus, it informs what position led to the feature being activated.
@@ A.15 Feature over-splitting in a mixture of gaussians toy model
We implemented this using $10^{5}$ samples, with a grid of 100 values for $\gamma$ in $[0,5]$ and 20 values for $k$ in $[0,2]$, over a grid of 100 values of $\lambda$ in $[0,20]$. We verify that the best values chosen for each $\lambda$ are not on the edges of the grid; the resulting curve of best total loss values versus $\lambda$ is shown in Figure 14 (blue), cut off at $\lambda\approx 7$, beyond which the selected SAE activates for $<2\%$ of examples in the components of the mixture.